OpenStudy (anonymous):

What are the x-intercepts of y=3x^2-17x+10?

OpenStudy (linn99123):

-3x2-17x-10=0 Two solutions were found : x = -5 x = -2/3 Step by step solution : Step 1 : Simplify -3x2-17x - 10 Pulling out like terms : 1.1 Pull out like factors : -3x2 - 17x - 10 = -1 • (3x2 + 17x + 10) Trying to factor by splitting the middle term 1.2 Factoring 3x2 + 17x + 10 The first term is, 3x2 its coefficient is 3 . The middle term is, +17x its coefficient is 17 . The last term, "the constant", is +10 Step-1 : Multiply the coefficient of the first term by the constant 3 • 10 = 30 Step-2 : Find two factors of 30 whose sum equals the coefficient of the middle term, which is 17 . -30 + -1 = -31 -15 + -2 = -17 -10 + -3 = -13 -6 + -5 = -11 -5 + -6 = -11 -3 + -10 = -13 -2 + -15 = -17 -1 + -30 = -31 1 + 30 = 31 2 + 15 = 17 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 2 and 15 3x2 + 2x + 15x + 10 Step-4 : Add up the first 2 terms, pulling out like factors : x • (3x+2) Add up the last 2 terms, pulling out common factors : 5 • (3x+2) Now add up the four terms of step 3 : (x+5) • (3x+2) Which is the desired factorization Equation at the end of step 1 : -1 • (3x + 2) • (x + 5) = 0 Step 2 : Solve -1•(3x+2)•(x+5) = 0 Theory - Roots of a product : 2.1 A product of several terms equals zero. When a product of two or more terms equals zero, then at least one of the terms must be zero. We shall now solve each term = 0 separately In other words, we are going to solve as many equations as there are terms in the product Any solution of term = 0 solves product = 0 as well. Equations which are never true : 2.2 Solve : -1 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.3 Solve : 3x+2 = 0 Subtract 2 from both sides of the equation : 3x = -2 Divide both sides of the equation by 3: x = -2/3 Solving a Single Variable Equation : 2.4 Solve : x+5 = 0 Subtract 5 from both sides of the equation : x = -5 Supplement : Solving Quadratic Equation Directly Solving 3x2+17x+10 = 0 directly Earlier we factored this polynomial by splitting the middle term. let us now solve the equation by Completing The Square and by using the Quadratic Formula Parabola, Finding the Vertex : 3.1 Find the Vertex of y = 3x2+17x+10 Parabolas have a highest or a lowest point called the Vertex . Our parabola opens up and accordingly has a lowest point (AKA absolute minimum) . We know this even before plotting "y" because the coefficient of the first term, 3 , is positive (greater than zero). Each parabola has a vertical line of symmetry that passes through its vertex. Because of this symmetry, the line of symmetry would, for example, pass through the midpoint of the two x -intercepts (roots or solutions) of the parabola. That is, if the parabola has indeed two real solutions. Parabolas can model many real life situations, such as the height above ground, of an object thrown upward, after some period of time. The vertex of the parabola can provide us with information, such as the maximum height that object, thrown upwards, can reach. For this reason we want to be able to find the coordinates of the vertex. For any parabola,Ax2+Bx+C,the x -coordinate of the vertex is given by -B/(2A) . In our case the x coordinate is -2.8333 Plugging into the parabola formula -2.8333 for x we can calculate the y -coordinate : y = 3.0 * -2.83 * -2.83 + 17.0 * -2.83 + 10.0 or y = -14.083 Parabola, Graphing Vertex and X-Intercepts : Root plot for : y = 3x2+17x+10 Axis of Symmetry (dashed) {x}={-2.83} Vertex at {x,y} = {-2.83,-14.08} x -Intercepts (Roots) : Root 1 at {x,y} = {-5.00, 0.00} Root 2 at {x,y} = {-0.67, 0.00} Solve Quadratic Equation by Completing The Square 3.2 Solving 3x2+17x+10 = 0 by Completing The Square . Divide both sides of the equation by 3 to have 1 as the coefficient of the first term : x2+(17/3)x+(10/3) = 0 Subtract 10/3 from both side of the equation : x2+(17/3)x = -10/3 Now the clever bit: Take the coefficient of x , which is 17/3 , divide by two, giving 17/6 , and finally square it giving 289/36 Add 289/36 to both sides of the equation : On the right hand side we have : -10/3 + 289/36 The common denominator of the two fractions is 36 Adding (-120/36)+(289/36) gives 169/36 So adding to both sides we finally get : x2+(17/3)x+(289/36) = 169/36 Adding 289/36 has completed the left hand side into a perfect square : x2+(17/3)x+(289/36) = (x+(17/6)) • (x+(17/6)) = (x+(17/6))2 Things which are equal to the same thing are also equal to one another. Since x2+(17/3)x+(289/36) = 169/36 and x2+(17/3)x+(289/36) = (x+(17/6))2 then, according to the law of transitivity, (x+(17/6))2 = 169/36 We'll refer to this Equation as Eq. #3.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (x+(17/6))2 is (x+(17/6))2/2 = (x+(17/6))1 = x+(17/6) Now, applying the Square Root Principle to Eq. #3.2.1 we get: x+(17/6) = √ 169/36 Subtract 17/6 from both sides to obtain: x = -17/6 + √ 169/36 Since a square root has two values, one positive and the other negative x2 + (17/3)x + (10/3) = 0 has two solutions: x = -17/6 + √ 169/36 or x = -17/6 - √ 169/36 Note that √ 169/36 can be written as √ 169 / √ 36 which is 13 / 6 Solve Quadratic Equation using the Quadratic Formula 3.3 Solving 3x2+17x+10 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 3 B = 17 C = 10 Accordingly, B2 - 4AC = 289 - 120 = 169 Applying the quadratic formula : -17 ± √ 169 x = —————— 6 Can √ 169 be simplified ? Yes! The prime factorization of 169 is 13•13 To be able to remove something from under the radical, there have to be 2 instances of it (because we are taking a square i.e. second root). √ 169 = √ 13•13 = ± 13 • √ 1 = ± 13 So now we are looking at: x = ( -17 ± 13) / 6 Two real solutions: x =(-17+√169)/6=(-17+13)/6= -0.667 or: x =(-17-√169)/6=(-17-13)/6= -5.000 Two solutions were found : x = -5 x = -2/3

OpenStudy (linn99123):

Two solutions were found : x = -5 x = -2/3

OpenStudy (anonymous):

this translates as "solve for \(x\) \[3x^2-17x+10=0\]

OpenStudy (anonymous):

ok can you tell me how to do this but not as much information?

OpenStudy (anonymous):

since by some miracle this factors as \[(3x-2)(x-5)=0\]the zeros are easy to find

OpenStudy (anonymous):

unfortunately the solutions given above are wrong

OpenStudy (anonymous):

\[3x-2=0\\ 3x=2\\ x=\frac{3}{2}\] is one zero

OpenStudy (anonymous):

\[x-5=0\\ x=5\] is the other one

OpenStudy (shadowlegendx):

@satellite73 Are we still friends c; @linn99123 Nice info on those subjects!

OpenStudy (linn99123):

lol ikr Shad xD

OpenStudy (anonymous):

yes of course not sure what all that stuff above is, looks like it could be right, but not necessary unfortunately the last step was wrong

OpenStudy (shadowlegendx):

@satellite73 Can we be more then friends? ;) @linn99123 Yeah :)